OpenStudy (anonymous):

x2-8x=33 using factor formula and quadratic formula. I need help

OpenStudy (anonymous):

Hint: x+3 is a factor.

OpenStudy (anonymous):

Solve it @Asuarez_7

OpenStudy (anonymous):

Thank you, can you help me on a other one?

OpenStudy (anonymous):

49^2-1=0

OpenStudy (anonymous):

wait,we haven't finish yet.. so what are values for x?

OpenStudy (anonymous):

*are the

OpenStudy (anonymous):

Is it 3?

OpenStudy (anonymous):

Wait no it's x= 11 and x= -3

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

One important issue should be mentioned at this point: Just as with linear equations, the solutions to quadratic equations may be verified by plugging them back into the original equation, and making sure that they work, that they result in a true statement. For the above example, we would do the following:\[checking~x=-3~inn~(x+3)(x-11)=0:\]\[([-3]+3)([-3]-11)=0\]\[(0)(-14)=0\]\[0=0\]\[checking~x=11~inn~(x+3)(x-11)=0:\]\[([11]+3)([11]-11)=0\]\[(14)(0)=0\]\[0=0\]So both solutions "check" and are thus verified as being correct.

OpenStudy (anonymous):

That was factorisation method,now let's use the quadratic formula to find the values for x

OpenStudy (anonymous):

Solving using Quadratic Formula \[x ^{2}-8x=33 \] \[Ax ^{2}+Bx+C=0\] \[x^{2}-8x-33=0\] a=1 b=-8 c=-33 \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] \[x=\frac{ -(-8) \pm \sqrt{8^{2}-4(1)(-33)} }{ 2(1) }\] \[x=\frac{ 8 \pm \sqrt{64+132} }{ 2 }\] \[x=\frac{ 8 \pm \sqrt{196} }{ 2 }\] \[x=\frac{ 8 \pm 14 }{ 2 }\] \[x _{1}=\frac{ 8 + 14 }{ 2 }\] \[x _{1}=\frac{ 22 }{ 2 } =11\] \[x _{2}=\frac{ 8 - 14 }{ 2 }\] \[x _{2}=-\frac{ 6 }{ 2 } = -3\] Hope this helped..

OpenStudy (anonymous):

Thanks it did but can you pls help me on this One using factoring and quadratic formula 49x^2-1=0

OpenStudy (anonymous):

Factorisation: Difference of two squares:\[x^2-a^2=(x-a)(x+a)\]\[49x^2-1=0\]\[7^2x^2-1^1=0\]\[(7x-1)(7x+1)\]\[7x-1=0,7x+1=0\]\[7x=1,7x=-1\]

OpenStudy (anonymous):

Then divide both sides by 7

OpenStudy (anonymous):

\[\frac{ 7x }{ 7 }=\frac{ 1 }{ 7 }\]\[\frac{ 7x }{ 7 }=-\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

so what are the values of x? @Asuarez_7

OpenStudy (anonymous):

Is it 7?

OpenStudy (anonymous):

no it is supposed to be\[x=\frac{ 1 }{ 7 }~and~x=-\frac{ 1 }{ 7 }\]

OpenStudy (anonymous):

One important issue should be mentioned at this point: Just as with linear equations, the solutions to quadratic equations may be verified by plugging them back into the original equation, and making sure that they work, that they result in a true statement. For the above example, we would do the following:\[checking~x=\frac{ 1 }{ 7 }~inn~(7x-1)(7x+1)=0:\]\[(7[\frac{ 1 }{ 7 }]-1)(7[\frac{ 1 }{ 7 }]+1)=0\]\[(0)(2)=0\]\[0=0\]\[checking~x=-\frac{ 1 }{ 7 }~inn~(7x-1)(7x+1)=0:\]\[(7[-\frac{ 1 }{ 7 }]-1)(7[-\frac{ 1 }{ 7 }]+1)=0\]\[(-2)(0)=0\]\[0=0\]So both solutions "check" and are thus verified as being correct.